OpenStudy (anonymous):

Jerry's haircut costs $15.00. He gave the barber a 10% tip. What was the total cost?

OpenStudy (anonymous):

$16.50

OpenStudy (anonymous):

15*0.10 = 1.5 + 15 = 16.50